# On Bessel’s Correction: Unbiased Sample Variance, the “Bariance,” and a Novel Runtime-Optimized Estimator

### v2

@@ 6 Discussion: Should We Just Divide by $n$?
In Rosenthal [1] argues that using $n$ instead of $n{-}1$ may lead to a smaller mean squared error (MSE) — especially when teaching or in practical settings.
He shows that while dividing by $n{-}1$ yields an unbiased estimator, this might come at the cost of higher variance. In some cases, a biased but lower-MSE estimator using $n$ is preferable:
@@
Numerical Example: Bias vs. MSE
Suppose we have $n=5$ observations drawn from a population with true variance $\sigma^{2}=10$. Then:
  • The biased estimator divides by $n=5$:  $S^{2}\approx\frac{4}{5}\cdot\sigma^{2}=8$
  • The unbiased estimator divides by $n-1=4$:  $\hat{S}^{2}=10$
Now compute Mean Squared Error (MSE):
$\text{MSE}(S^{2})$ $=\underbrace{\left(\operatorname{\mathbb{E}}[S^{2}]-\sigma^{2}\right)^{2}}_{\text{Bias}^{2}}+\underbrace{\operatorname{\mathrm{Var}}(S^{2})}_{\text{Variance}}$
$\text{MSE}(\hat{S}^{2})$ $=\underbrace{0^{2}}_{\text{= Unbiased (Bessel's correction)}}+\operatorname{\mathrm{Var}}(\hat{S}^{2})$
It turns out (and Rosenthal notes this explicitly) that: - $\operatorname{\mathrm{Var}}(S^{2})<\operatorname{\mathrm{Var}}(\hat{S}^{2})$ So in some cases, even though $S^{2}$ is biased, its total MSE is still smaller!
@@ 7 A Simulation Study: Bias2, Variance, and MSE Across Denominator Values
Table 1: Empirical results averaged from 100,000 simulations for variance estimator using $n=5$ and $a\in[3.5,8.5]$ in 0.5 increments. The bold rows highlight $a\in$ $n-1$, $n$ and $n+1$, respectively.
a (denominator) Bias2 Variance MSE
3.5 2.1044 65.8077 67.9122
4.0 0.0004 50.3840 50.3844
4.5 1.1967 39.8096 41.0063
5.0 3.9384 32.2458 36.1842
5.5 7.3615 26.6494 34.0109
6.0 11.0254 22.3929 33.4183
6.5 14.7015 19.0803 33.7819
7.0 18.2728 16.4519 34.7247
7.5 21.6817 14.3315 36.0131
8.0 24.9034 12.5960 37.4994
8.5 27.9314 11.1577 39.0891
33333.53.53.53.544444.54.54.54.555555.55.55.55.566666.56.56.56.577777.57.57.57.588888.58.58.58.59999001010101020202020303030304040404050505050Denominator $a$ValueMSE, Bias2, and Variance of Variance EstimatorMSEBias2VarianceDivide by $n{-}1$Divide by $n$
Figure 5: Empirical MSE, Bias2, and Variance of the sample variance estimator for $a\in[3.5,8.5]$ and $n=5$. Minimum MSE occurs between $a=5.5$ and $a=6.5$. ($16=n+$ 1 in the case of $\mathcal{N}$)
8 Computational Complexity of Variance Bariance Estimators and Optimization
Let $X=\{X_{1},X_{2},\dots,X_{n}\}\subset\mathbb{R}$ be a sample of size $n$. Define:
  • Biased sample variance:
    $S^{2}:=\frac{1}{n}\sum_{i=1}^{n}(X_{i}-\bar{X})^{2}$
  • Unbiased sample variance (Bessel corrected):
    $\hat{S}^{2}:=\frac{1}{n-1}\sum_{i=1}^{n}(X_{i}-\bar{X})^{2}$
  • Bariance (pairwise variance):
    $\text{Bariance}(X):=\frac{1}{n(n-1)}\sum_{i\neq j}(X_{i}-X_{j})^{2}$
@@ 8.1 Computational Complexity Comparison with Numerical Illustration
$\bar{X}=6,\quad S^{2}=\frac{1}{5}\sum(X_{i}-6)^{2}=\frac{40}{5}=8$
@@
$\sum_{i<j}(X_{i}-X_{j})^{2}=200,\quad\text{(10 unordered pairs)}\Rightarrow\text{Bariance}=\frac{2\cdot 200}{20}=20$
@@
$\sum X_{i}=30,\quad\sum X_{i}^{2}=220$
$\text{Bariance}=\frac{2\cdot 5}{5\cdot 4}\cdot 220-\frac{2}{5\cdot 4}\cdot 900=110-90=20$
Thus, all estimators yield consistent results, but the number of operations differs significantly with growing $n$.
While the pairwise form of the bariance appears quadratic, algebraic reduction allows it to be computed in linear time, just like classical variance. This makes it a viable alternative even in large-scale statistical computations.
8.2 Empirical Runtime
@@
8.2.1 Parameters of the Normal-Based Simulation
@@
  • Hardware environment: CPU timing measured in Python on a standard workstation
@@
Table 3: Empirical runtime (in seconds) for 1000 simulations per estimator at different sample sizes
@@
10101010202020203030303040404040505050506060606070707070808080809090909010010010010000222244446666Sample Size $n$Runtime (seconds)Empirical Runtime for 1000 Simulations per EstimatorBiasedUnbiasedNaive BarianceOptimized Bariance
Figure 6: Empirical runtime comparison of variance and bariance estimators over 1000 simulations per sample size.
8.2.2 Gamma-Distributed Data
@@
  • Hardware environment: Standard workstation with vectorized Python implementation
Table 4: Empirical runtime (in seconds) for 500 simulations per estimator using Gamma-distributed data
@@
8.2.3 Highly Dispersed Gamma-Distributed Data
@@
  • Hardware environment: Standard workstation with vectorized Python implementation
Table 5: Empirical runtime (in seconds) for 1000 simulations using a highly dispersed Gamma distribution
@@ 9 Conclusion
Bessel’s correction is a foundational concept that ensures unbiased estimates of variance. We explored its necessity through algebraic, geometric, and pairwise differences reasoning (“Bariance“), building both intuition and understanding. Additionally, we considered a pedagocial and practical perspective, such as Rosenthal’s $MSE$-based view for estimating variance.
Although the unbiased estimator is mathematically correct in expectation, the biased version can sometimes be more intuitive, and, in certain contexts, is statistically preferable in most cases – for many sampling distributions. Furthermore, the empirical results reveal a faster runtime in our simulation example using the average pairwise differences definition as unbiased variance estimator when using the algebraically optimized definition with scalar sums. Further research could focus on experiments using unbiased or shrunken runtime-optimized covariance estimators with the application of speeding up construction of variance-covariance matrices.

### v5

@@ 6 Discussion: Should We Just Divide by $n$?
Rosenthal [8] argues that using $n$ instead of $n-1$ may lead to a smaller mean squared error (MSE) — especially when teaching or in practical settings.
He shows that while dividing by $n-1$ yields an unbiased estimator, this may come at the cost of increased variance. In some cases, a biased but lower-MSE estimator using $n$ is preferable:
@@
From a theoretical perspective, unbiasedness ensures that the expected value of the estimator exactly matches the true population variance. However, unbiasedness alone does not guarantee minimal estimation error in finite samples. In fact, particularly when sample sizes are small, the variance of the unbiased estimator can be relatively large, which may lead to unstable or inefficient estimates. Allowing a small bias can reduce this variance enough to produce a lower overall mean squared error, which combines both bias and variance in a single measure of estimator quality [1, 10].
Practically, this trade-off becomes important in many applied contexts, such as when the variance estimate is an intermediate quantity used for further modeling or prediction. Here, reducing total estimation error takes precedence over strict unbiasedness. Additionally, computational simplicity and pedagogical clarity sometimes favor the $n$-denominator estimator, which is easier to understand and implement, especially in introductory settings.
To illustrate, consider the case of $n=5$ observations drawn from a population with true variance $\sigma^{2}=10$. The biased estimator, which divides by $n=5$, underestimates the variance as $8$, while the unbiased estimator, dividing by $n-1=4$, correctly yields $10$. However, when mean squared error is calculated, which accounts for both bias and variance, the biased estimator can have smaller total error due to its reduced variance. Rosenthal explicitly notes that the variance of the biased estimator is often less than that of the unbiased one, which explains this result.
This nuanced understanding clarifies why and when the classical insistence on unbiasedness may be relaxed in favor of better finite-sample performance and practical utility. Furthermore, more generalized estimators with a denominator parameter $a>0$, defined as
$\hat{\sigma}^{2}_{a}:=\frac{1}{a}\sum_{i=1}^{n}(X_{i}-\bar{X})^{2},$
can be analyzed for optimality in terms of mean squared error. As derived in Appendix B, the MSE-minimizing denominator is approximately $a^{*}\approx n+1$, suggesting that neither the classical unbiased divisor $n-1$ nor the biased $n$ divisor are necessarily optimal from an MSE standpoint.
In sum, relaxing unbiasedness for variance estimation is a principled and context-dependent choice motivated by the bias-variance trade-off, sample size considerations, and the ultimate goals of estimation. This perspective complements classical theory and better reflects the realities of applied statistical practice.
@@ 7 A Simulation Study: Bias2, Variance, and MSE Across Denominator Values
Table 1: Empirical Bias2, Variance, and MSE with 95% bootstrapped confidence intervals (200 resamples, seed=42). Bold rows indicate $a=n-1$, $n$, and $n+1$. Hardware: 1GB RAM, Python 3.11. For a theoretical derivation of the MSE-minimized variance estimator, see Appendix B.
$a$ Bias2 [CI] Variance [CI] MSE [CI]
3.5 1.98 [1.97, 2.03] 61.99 [61.85, 62.19] 63.96 [63.83, 64.21]
4.0 0.00 [0.00, 0.01] 47.46 [47.26, 47.51] 47.46 [47.26, 47.52]
4.5 1.27 [1.26, 1.30] 37.50 [37.34, 37.57] 38.77 [38.63, 38.84]
5.0 4.06 [4.02, 4.08] 30.37 [30.29, 30.47] 34.44 [34.36, 34.50]
5.5 7.52 [7.48, 7.54] 25.10 [25.07, 25.20] 32.62 [32.59, 32.70]
6.0 11.20 [11.16, 11.24] 21.09 [21.05, 21.17] 32.29 [32.27, 32.36]
6.5 14.89 [14.81, 14.90] 17.97 [17.92, 18.03] 32.86 [32.79, 32.87]
7.0 18.46 [18.46, 18.56] 15.50 [15.42, 15.51] 33.96 [33.94, 34.01]
7.5 21.88 [21.85, 21.94] 13.50 [13.48, 13.55] 35.37 [35.37, 35.44]
8.0 25.10 [25.06, 25.17] 11.86 [11.82, 11.89] 36.96 [36.93, 37.01]
8.5 28.13 [28.09, 28.20] 10.51 [10.45, 10.52] 38.64 [38.59, 38.67]
33333.53.53.53.544444.54.54.54.555555.55.55.55.566666.56.56.56.577777.57.57.57.588888.58.58.58.59999005555101010101515151520202020252525253030303035353535404040404545454550505050Denominator $a$ValueMSE, Bias2, and Variance of Variance EstimatorMSEBias2VarianceDivide by $n{-}1$Divide by $n$Divide by $n{+}1$
Figure 5: Empirical MSE, Bias2, and Variance of the sample variance estimator for $a\in[3.5,8.5]$ and $n=5$ using 10,000 simulations and 200 bootstraps. Minimum MSE occurs between $a=5.5$ and $a=6.5$. For a theoretical derivation of the MSE-minimized variance estimator, see Appendix B.
8 On Computational Complexity of Variance Bariance Estimators and Optimization
Let $X:=\{X_{1},X_{2},\dots,X_{n}\}\subset\mathbb{R}$ be a sample of size $n$.
@@ 8.1 Computational Complexity Comparison with Numerical Illustration
$S^{2}=\frac{1}{5}\sum(X_{i}-\bar{X})^{2}=\frac{1}{5}[(1-5)^{2}+(3-5)^{2}+(5-5)^{2}+(7-5)^{2}+(9-5)^{2}]=\frac{1}{5}[16+4+0+4+16]=\frac{40}{5}=8$
@@
$\sum_{i<j}(X_{i}-X_{j})^{2}=(3-1)^{2}+(5-1)^{2}+(7-1)^{2}+(9-1)^{2}+(5-3)^{2}+(7-3)^{2}+(9-3)^{2}+(7-5)^{2}+(9-5)^{2}+(9-7)^{2}$
$=4+16+36+64+4+16+36+4+16+4=200$
$\text{{Bariance}}=\frac{2\cdot 200}{5\cdot 4}=\frac{400}{20}=20$
@@
$\sum X_{i}=1+3+5+7+9=25,\quad\sum X_{i}^{2}=1^{2}+3^{2}+5^{2}+7^{2}+9^{2}=165$
$\text{{Bariance}}=\frac{2n}{n(n-1)}\sum X_{i}^{2}-\frac{2}{n(n-1)}\left(\sum X_{i}\right)^{2}=\frac{2\cdot 5}{20}\cdot 165-\frac{2}{20}\cdot 625=\frac{1650}{20}-\frac{1250}{20}=82.5-62.5=20$
Thus, all estimators yield consistent results, confirming their correctness. However, their computational complexity differs:
  • Biased/Unbiased Variance: $\mathcal{O}(n)$
  • Naïve Bariance: $\mathcal{O}(n^{2})$
  • Optimized Bariance: $\mathcal{O}(n)$
The optimized Bariance offers the same result as the naïve form but with significantly reduced computational cost, making it efficient for large-scale applications.
9 On Empirical Runtime of Bariance Estimators
@@
9.1 Execution Environment
All Python code was executed in a virtualized Python 3.11.8 environment on a Linux system (kernel version 4.4.0) with x86_64 architecture. The processor was identified as unknown, featuring 32 logical cores and 32 physical cores. The system reported a BogoMIPS value of 2593.91.
Available memory was limited to 1.07 GB of RAM, with no swap configured. Supported CPU instruction sets included AVX, AVX2, AVX-512, and FMA, as listed in /proc/cpuinfo.
All timing measurements were performed using time.perf_counter to capture high-resolution wall-clock time in single-threaded execution mode. Each estimator was evaluated across $\tau$ = 20 independent trials per sample size.
9.2 Normal-Distributed Data
@@
Table 3: Empirical runtime (in seconds) for 1,000 simulations per estimator across different sample sizes ($n$). Time measured in wall-clock seconds. Bold values indicate the fastest method for each row.
@@
9.3 Gamma-Distributed Data
@@
Table 4: Empirical runtime (in seconds) for 500 simulations per estimator using Gamma-distributed data. Time measured in wall-clock seconds. Bold values highlight the fastest method at each sample size $n$.
@@
9.4 Highly Dispersed Gamma-Distributed Data
@@
Table 5: Empirical runtime (in seconds) for 1,000 simulations per estimator using highly dispersed Gamma-distributed data. Time measured in wall-clock seconds. Bold values indicate the fastest method for each sample size $n$.
@@
9.5 Robustness: Statistical Analysis of Empirical Runtime
For Robustness testing in an alternative execution environment see Appendix C and the replication package  [^1] https://github.com/felix-reichel/BarianceVariance_Reproduction_Repo_Robustness_Supplements.
9.5.1 Execution Environment
Available memory was limited to 1.07 GB of RAM, with no swap configured. Supported CPU instruction sets included AVX, AVX2, AVX-512, and FMA, as listed in /proc/cpuinfo.
All timing measurements were performed using time.perf_counter to capture high-resolution wall-clock time in single-threaded execution mode. Each estimator was evaluated across $\tau$ = 20 independent trials per sample size.
9.5.2 Simulation Setup
All estimators were tested on synthetic data generated from a standard normal distribution $\mathcal{N}(0,1)$. The following estimators were implemented:
  • Biased and Unbiased Variance (Looped and Vectorized)
  • Optimized Bariance (Looped and Vectorized)
Sample sizes ranged from 100 to 4800 in even steps for a total of 48 tests, with an additional focused study at 16 evenly spaced sizes for detailed visualization. See Figure 7.
9.5.3 Main Estimator Comparison
Figure 6 shows kernel density estimates of runtime differences between the Unbiased Vectorized and Optimized Bariance Vectorized estimators across 16 sample sizes. The densities summarize performance variability over multiple runs.
Figure 6: Kernel density estimates of runtime differences between Unbiased Sample Variance Vectorized and Optimized Bariance Vectorized estimators. Positive values indicate slower performance of the unbiased estimator.
9.5.4 Kernel Density Comparison: 3x16 Grid
Figure 7 presents a 3x16 grid of kernel density estimates for runtime differences between the same two estimators across 48 sample sizes. Each panel shows the distribution for a distinct sample size $n$.
Figure 7: Kernel density estimates of runtime differences for Unbiased Sample Variance Vectorized and Optimized Bariance Vectorized estimators across 48 sample sizes.
9.5.5 All Estimators: Comparative Runtime Analysis
Figure 8 displays kernel density estimates of runtime for all estimators over 16 sample sizes. The two primary estimators are emphasized in color, while others appear in grayscale for contrast.
Figure 8: Kernel density estimates of runtime for all estimators across 16 sample sizes. Unbiased Sample Variance Vectorized and Optimized Bariance Vectorized are highlighted; grayscale lines represent additional looped and vectorized variants.
@@ 10 Conclusion
Bessel’s correction is a foundational concept that ensures unbiased estimates of variance. We explored its necessity through algebraic, geometric, and pairwise differences reasoning (now formalized as the Bariance construct), building both intuition and understanding. Additionally, we considered a pedagogical and practical perspective, such as Rosenthal’s $MSE$-based view for estimating variance [8].
Although the unbiased estimator is mathematically correct in expectation, the biased version can sometimes be more intuitive and, in certain contexts, statistically preferable across various sampling distributions. This aligns with insights from modern treatments of mathematical statistics, which often emphasize the trade-off between bias and variance in estimator performance [1, 10]. Furthermore, empirical results revealed a faster runtime in our simulation example using the average pairwise differences definition as an unbiased variance estimator—referred to as the Bariance estimator—particularly when employing the algebraically optimized formula using scalar sums.
To sum up, the main finding—the run-time optimized estimator for the Bariance formula—was a coincidental yet significant observation: that is, the unbiased estimator can be computed in linear time and statistically outperforms the conventional unbiased sample variance estimator in all tested empirical runtime performance scenarios. Naturally, many other estimators exist for sample variance, including those designed to trade off bias for computational gains. A complexity theorist or mathematician could potentially derive theoretical bounds on the time complexity of such estimators.
Beyond its computational efficiency, however, the Bariance measure may also offer substantive benefits in applied settings where deviation from a central mean is either unstable, undefined, or conceptually inappropriate. In fields such as genomics, network analysis, robust statistics, or ordinal survey research, dispersion may be more meaningfully characterized by average pairwise differences rather than deviations from a global average. Moreover, distance-based methods like clustering, energy statistics, and nonparametric ANOVA can all benefit from the geometric and symmetry-preserving properties of Bariance, particularly in high-dimensional or irregularly structured data where the mean offers little interpretive value. These contexts highlight how the pairwise construction of Bariance is not only computationally attractive but also methodologically appropriate.
Thus, the optimized Bariance formula stands as a viable alternative with promising practical implications for real-time multivariate big data applications, including forecasting (especially with shrunken variance-covariance estimators), computational biology, chemistry, finance, and big-data streaming applications (such as online learning) where unbiased and scalable variance estimation is essential.
@@ Conclusion
Appendix B Derivation of MSE-Optimal Denominator for Variance Estimator
Proof.
We consider family of $a$-based denominator estimators of sample variance:
$\hat{\sigma}^{2}_{a}:=\frac{1}{a}\sum_{i=1}^{n}(X_{i}-\bar{X})^{2}$
Assume $X_{i}\sim\mathcal{N}(\mu,\sigma^{2})$ i.i.d.
We start with the Bias of the estimator:
$\mathbb{E}[\hat{\sigma}^{2}_{a}]=\frac{n-1}{a}\sigma^{2}\quad\Rightarrow\quad\text{Bias}=\mathbb{E}[\hat{\sigma}^{2}_{a}]-\sigma^{2}=\left(\frac{n-1-a}{a}\right)\sigma^{2}$
$\Rightarrow\text{Bias}^{2}=\left(\frac{n-1-a}{a}\right)^{2}\sigma^{4}$
Furthermore, for the Variance it is known that:
$\operatorname{Var}\left(\sum_{i=1}^{n}(X_{i}-\bar{X})^{2}\right)=2(n-1)\sigma^{4}$
Thus:
$\operatorname{Var}[\hat{\sigma}^{2}_{a}]=\operatorname{Var}\left(\frac{1}{a}\sum_{i=1}^{n}(X_{i}-\bar{X})^{2}\right)=\frac{1}{a^{2}}\cdot 2(n-1)\sigma^{4}=\frac{2\sigma^{4}}{a^{2}}(n-1)$
Computing the Mean-Squared Error (MSE):
$\text{MSE}(a)=\text{Bias}^{2}+\text{Var}=\left(\frac{n-1-a}{a}\right)^{2}\sigma^{4}+\frac{2(n-1)}{a^{2}}\sigma^{4}$
Factoring out $\sigma^{4}$:
$\text{MSE}(a)=\sigma^{4}\left[\left(\frac{n-1-a}{a}\right)^{2}+\frac{2(n-1)}{a^{2}}\right]$
Let:
$f(a):=\left(\frac{n-1-a}{a}\right)^{2}+\frac{2(n-1)}{a^{2}}=\frac{(n-1-a)^{2}+2(n-1)}{a^{2}}$
We seek to minimize $f(a)$ over $a>0$.
Minimization
Let:
$f(a):=\frac{u(a)}{v(a)}\quad\text{with}\quad u(a):=(n-1-a)^{2}+2(n-1),\quad v(a):=a^{2}$
Compute derivatives:
$u^{\prime}(a)=-2(n-1-a),\qquad v^{\prime}(a)=2a$
$f^{\prime}(a)=\frac{u^{\prime}(a)v(a)-u(a)v^{\prime}(a)}{v(a)^{2}}=\frac{-2(n-1-a)a^{2}-2a\left[(n-1-a)^{2}+2(n-1)\right]}{a^{4}}$
Set numerator = 0:
$-2(n-1-a)a^{2}-2a\left[(n-1-a)^{2}+2(n-1)\right]=0$
This equation is non-linear in $a$; solving analytically is messy, but plugging into a symbolic solver yields:
$\boxed{a^{*}\approx n+1}$
which thus is the MSE-minimizing choice of denominator $a$ and is approximately as in previously shown simulations and [8].
Appendix C Robustness Check of Empirical Runtime in Java SE
To verify the consistency of results across platforms, we implemented both the unbiased and optimized Bariance estimators in Java 21.0.1 on a system configured as follows: Mac OS X 13.0 operating system; aarch64 architecture; 10 cores available (single-threaded execution); and 4 GB of JVM-reported maximum memory.
Runtime Results
C.1 Normal-Distributed Data
Each estimator was benchmarked over $\tau$ = 100 trials for each sample size. Figure 9 displays the mean runtime with 95% confidence intervals for both estimators.
Figure 9: Candlestick plots (mean $\pm$ 1.96 standard errors) of Java runtime estimates for unbiased and optimized Bariance estimators.
Kernel Density Estimates
Figure 10 illustrates the runtime distribution for each estimator across selected sample sizes, visualized via kernel density estimates.
Figure 10: Density plots of runtime for unbiased sample variance and optimized Bariance estimators in Java, stratified by sample size.
C.2 Gamma-Distributed Data
Table 7: OLS Regression of Runtime on Estimator and Sample Size
Coefficient Std. Error
Intercept 698.33*** (25.85)
Estimator: Optimized Bariance -909.21*** (29.46)
Estimator: Biased -694.43*** (28.91)
Estimator: Population Variance 1457.29*** (25.56)
Estimator: Unbiased -690.49*** (28.98)
Sample Size = 500 229.27*** (27.55)
Sample Size = 1000 534.47*** (29.38)
Sample Size = 2000 800.55*** (31.78)
Sample Size = 3000 1656.77*** (31.82)
Sample Size = 5000 2221.24*** (32.95)
R-squared 0.455
Observations 22,906
Notes: Dependent variable is runtime in nanoseconds. The model was estimated via ordinary least squares with fixed effects for estimator and sample size. Runtime was measured using Java’s System.nanoTime. Each estimator was run over $\tau$ = 1,000 trials per sample size with data drawn from a $\Gamma(2.0,2.0)$ distribution, seeded at 42. All implementations were validated with a synthetic test suite.
Figure 11: Estimated mean runtimes (nanoseconds) for five variance estimators across six sample sizes. Each point reflects the mean runtime estimated from $\tau$ = 1,000 Monte Carlo trials using Java’s System.nanoTime function. Error bars represent 95% confidence intervals obtained from a fixed effects ordinary least squares regression on sample size and estimator. Data were generated from a $\Gamma(2.0,2.0)$ distribution with fixed seed (42). Classical estimators include the biased and unbiased sample variance and the population variance. Bariance-based methods refer to the optimized and alternative scalar formulations. Naïve pairwise estimators were excluded due to high computational cost. See Table 7 for regression coefficients.
Kernel Density Estimates
Figure 12 illustrates the runtime distribution for each estimator across selected sample sizes, visualized via kernel density estimates.
Figure 12: Kernel density estimates of runtime (in nanoseconds) for the unbiased sample variance and optimized Bariance variance estimators across nine increasing sample sizes (from $n=100$ to $n=500{,}000$). Each panel is based on $\tau$ = 1,000 trials using data sampled from a $\Gamma(2.0,2.0)$ distribution. Vertical dashed lines indicate the mean runtime per estimator and sample size.